OpenStudy (anonymous):

x + 1 - 2 square root x + 4 = 0

OpenStudy (anonymous):

\[x+1-2\sqrt{x+4}=0\] is this what the equation is supposed to look like?

OpenStudy (anonymous):

yes, thank you

OpenStudy (anonymous):

no problem, so you have to isolate the x, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so -2 sr x + 4 = -x -1

OpenStudy (anonymous):

ok, so the x+1 is attached? because there are no parenthesis around them so i dont believe they are attached to each other, which means you can subtract just thee one so you will get: x-2srx+4=-1

OpenStudy (anonymous):

Then is it x squared + 16 x - 32 = -1?

OpenStudy (anonymous):

add \(2\sqrt{x+4}\) to both sides and start with \[x+1=2\sqrt{x+4}\] then square both sides

OpenStudy (anonymous):

Ahhh, ok I think I see it now and in the end x = 5?

OpenStudy (anonymous):

5 works for sure you might get two answers though, maybe not

OpenStudy (anonymous):

Yes, of course. You rock! Thanks!

OpenStudy (anonymous):

oh no only one solution, just 5

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Ok, phew. I was just checking to be sure lol

OpenStudy (anonymous):

good work!

OpenStudy (anonymous):

:)